OpenStudy (anonymous):

An object falls from rest on a high tower and takes 5.0 s to hit the ground. Calculate the object’s position from the top of the tower at 1.0 s intervals. Make a position-time graph for the object’s motion. In your response, show what you are given, the equation that you used, any algebra required, a table of data, and your graph. g = 9.8 m/s^2

OpenStudy (aum):

\[ s = ut + \frac 12 gt^2 \\ u = 0 \\ s = \frac 12 * 9.8 * t^2 = 4.9t^2 \]Put t = 1, 2, 3, .... and calculate s.